OpenStudy (anonymous):

complete the table below and write a rule relating x and y x -3,-2,-1,0,1,2,3,4 y -7, ?, ?, 2, 5,?,?,14

jimthompson5910 (jim_thompson5910):

try to find the equation of the line through (0,2) and (1,5)

OpenStudy (anonymous):

x=3/1+2 ?

jimthompson5910 (jim_thompson5910):

very close

jimthompson5910 (jim_thompson5910):

but no

jimthompson5910 (jim_thompson5910):

remember that y = mx+b is the general slope intercept form

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

what did you get for the slope

OpenStudy (anonymous):

1/3

jimthompson5910 (jim_thompson5910):

incorrect

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

3/1?

jimthompson5910 (jim_thompson5910):

yes or 3

jimthompson5910 (jim_thompson5910):

y-intercept is what?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

i'm not sure

jimthompson5910 (jim_thompson5910):

the y-intercept is the y value of the point when x = 0

OpenStudy (anonymous):

ooh ok, then 2?

jimthompson5910 (jim_thompson5910):

so b = 2

jimthompson5910 (jim_thompson5910):

m = 3, b = 2

jimthompson5910 (jim_thompson5910):

the equation y = mx+b turns into y = 3x+2

jimthompson5910 (jim_thompson5910):

you can now determine the other y values

OpenStudy (anonymous):

oooh thank you I understand now

jimthompson5910 (jim_thompson5910):

np